(3-y)/x^2  dx+(y^2- 2x)/(xy^2 ) dy=0        y(-1)=2


1
Expert's answer
2020-10-01T14:32:43-0400

The given differential equation is "\frac{3-y}{x^2}dx \space +\frac{y^2-2x}{xy^2}dy=0" .


The given differential equation is of first order and first degree and can be expressed in the form of

"Mdx \space + Ndy =0" ,where M and N are functions of x and y .


compairing with standard form we get "M = \frac{3-y}{x^2} \space and \space N = \frac{y^2-2x}{xy^2}"


"Now \space \frac{\delta M}{\delta y }= -1/x^2" ( partially differentiating M with respect to y keeping x as constant)

"and \space \frac{\delta N}{\delta x }= \frac {\delta }{\delta x}(y^2/xy^2 - 2x/xy^2)\\or, \frac{\delta N}{\delta x }= \frac {\delta }{\delta x}(1/x - 2/y^2)\\or, \frac{\delta N}{\delta x }=-1/x^2"( partially differentiating N with respect to x keeping y as constant)


"As \space \space \frac{\delta M}{\delta y }= \frac{\delta N}{\delta x }"

"\therefore" The given differential equation is exact.

If the equation is exact ,then for some function "u(x,y)" we have

"Mdx \space + Ndy =" a perfect differential of "u = du."


Now the general solution can be obtained by the following rules :

  1. integrate M dx as if y is constant
  2. integrate N dy as if x is constant
  3. delete the term from 2 which is already obtained in 1.
  4. the sum of these integrals will give a function "u(x,y)" such that "du = M dx + N dy."


1) "\int ( \frac{3-y}{x^2}) dx = (3-y) \int \frac{1}{x^2}dx"

"= (3-y)\frac{(-1)}{x}"

"= y/x \space - 3/x."

2) "\int \frac{y^2-2x}{xy^2} dy = \int (y^2/xy^2 - 2x/xy^2) dy"

"=\int (1/x \space - 2/y^2) dy"

"= \int(1/x) dy \space - 2 \int(1/y^2)dy"

"= y/x \space + 2/y."

3) "u = \frac{y}{x} - \frac{3}{x} + \frac{2}{y}" (deleting the term (y/x) already obtained in 1.)

4) The required primitive is

"\frac{y}{x} - \frac{3}{x} + \frac{2}{y} = C"


"Now \space when \space x=-1, y = 2."

Therefore "\frac{2}{-1} - \frac{3}{(-1)} + \frac{2}{2} = C"

"\therefore C = -2 +3 +1 =2"


The final solution is

"\frac{y}{x} - \frac{3}{x} + \frac{2}{y} = 2" (ANSWER)